OpenStudy (anonymous):

How do I write -2+2i in polar form??

OpenStudy (shinalcantara):

-2 + 2i is in the form x + yi getting the value of r you'll have: \[r = \sqrt{(-2)^2 +(2)^2}\] \[r = \sqrt{8}\] \[r = 2\sqrt{2}\] for the angle you'll have: \[\tan \theta = \frac{ y }{ x }\] \[\tan \theta = \frac{ 2 }{ -2 }\] note that in this case x is negative and y is positive therefore we can say that the angle is in the second quadrant \[\theta = \tan^{-1} (-1)\] \[\theta = -45^o\] since we know that it's on the second quadrant then we'll have the angle as 180-45=135 therefore you'll have the polar form as \[r \angle \theta\] \[2\sqrt{2} \angle 135^o\]